# On Marginal Stability in Low Temperature Spherical Spin Glasses

### v2

@@ 2.2 No Upward Outliers in Even Models
${\mathrm{GS}}_{2,\varepsilon}(H_{N})=\sup_{\begin{subarray}{c}{\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime}\in{\mathcal{S}}_{N}\\ R({\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime})=1-\varepsilon\end{subarray}}\frac{H_{N}({\boldsymbol{\sigma}})+H_{N}({\boldsymbol{\sigma}}^{\prime})}{N}.$
@@
The multi-replica Parisi formula is given by a similar variational problem as in Proposition 1.1. We choose suitable interpolation parameters to upper bound ${\mathrm{GS}}_{2,\varepsilon}(\xi)$ for small $\varepsilon$, and thus deduce non-existence of outlier eigenvalues at approximate ground states. The general formula replaces $L$ by a $2\times 2$ positive semi-definite matrix $\boldsymbol{L}$, and $\zeta$ by a cumulative distribution function on a monotone path of $2\times 2$ matrices in the positive semi-definite order. The path of matrices is encapsulated by an function $\Phi:[0,2]\to{\mathbb{R}}^{2\times 2}$ with $\Phi(0)=0,\Phi(2)=Q$ and ${\sf Tr}(\Phi(t))=t$ for all $t$, with $\Phi(t)-\Phi(s)$ positive semi-definite for all $t\geq s$. Meanwhile $\alpha:[0,1]\to{\mathbb{R}}_{+}$ is the associated cumulative distribution function, which can be viewed as a positive measure on the range of $\Phi$. For us, the relevant specialization is as follows. Define the $2\times 2$ matrices:
@@
We take as given $(\zeta,L)=\operatorname*{\rm arg\,min}_{(\zeta,L)\in{\mathcal{K}}}{\mathcal{Q}}(\zeta,L;\xi)$. With $\ell=\xi^{\prime\prime}(1)^{-1/2}$, their two-replica generalizations are:
$\boldsymbol{L}=\frac{LJ_{+}+\ell J_{-}}{2};\quad\quad\alpha(t)=\zeta(t/2)/2;\quad\quad\Phi(t)=\begin{cases}\frac{tJ_{+}}{2},\quad t\in[0,2-\varepsilon];\\ \frac{(2-\varepsilon)J_{+}}{2}+\frac{t-(2-\varepsilon)J_{-}}{2},\quad t\in[2-\varepsilon,2].\end{cases}$
@@
$2\,{\mathrm{GS}}_{2,\varepsilon}(\xi)$ $\leq\langle\xi^{\prime}(Q),\boldsymbol{L}\rangle-\int_{0}^{2}\Big{\langle}\xi^{\prime\prime}(\Phi(t))\odot\Phi^{\prime}(t),\int_{0}^{t}\alpha(s)\Phi^{\prime}(s)\,{\rm d}s\rangle\,{\rm d}t\Big{\rangle}$
@@
We evaluate the upper bound for $2{\mathrm{GS}}_{2,\varepsilon}(\xi)$ in Proposition 2.4 to first order in $\varepsilon$. The first term is
$L\big{(}\xi^{\prime}(1)+(1-\varepsilon)\xi^{\prime}(1)\big{)}=2L\xi^{\prime}(1)-L\xi^{\prime\prime}(1)\varepsilon+\ell\xi^{\prime\prime}(1)\varepsilon+O(\varepsilon^{2}).$
@@
$\int\limits_{0}^{2-\varepsilon}\left\langle\frac{\xi^{\prime\prime}(t/2)J_{+}}{2},J_{+}\int_{0}^{t}\zeta(s/2)/4\leavevmode\nobreak\ {\rm d}s\right\rangle{\rm d}t$ $=2\int\limits_{0}^{1-\frac{\varepsilon}{2}}\xi^{\prime\prime}(r)\left(\int_{0}^{v}\zeta(u)\,{\rm d}u\right){\rm d}v$
$=2\int_{0}^{1}\xi^{\prime\prime}(r)\left(\int_{0}^{v}\zeta(u)\,{\rm d}u\right){\rm d}v-\left(\xi^{\prime\prime}(1)\int_{0}^{1}\zeta(u)\,{\rm d}u\right)\varepsilon+O(\varepsilon^{2}).$
@@
$\Big{\langle}O(\varepsilon)J_{+}+\Theta(1)J_{-},\leavevmode\nobreak\ \Theta(1)J_{+}+O(\varepsilon)J_{-}\Big{\rangle}=O(\varepsilon).$
For the third term, we have $\Phi^{\prime}(t)=J_{+}/2$ on $t\in[0,2-\varepsilon]$. Note that
$(AJ_{+}+BJ_{-})^{-1}=(A^{-1}J_{+}+B^{-1}J_{-})/4,\quad\forall A,B\neq 0.$ (2.5)
Hence we only need to track the $J_{+}$ component in the matrix $\boldsymbol{L}-\int_{0}^{t}\alpha(s)\Phi^{\prime}(s){\rm d}s$ to determine the integral contribution on $[0,2-\varepsilon]$. This yields:
$\int_{0}^{2-\varepsilon}\left(\frac{L}{2}-\frac{1}{2}\int_{0}^{t/2}\zeta(u){\rm d}u\right)^{-1}\frac{1}{2}\leavevmode\nobreak\ {\rm d}t$ $=\int_{0}^{2-\varepsilon}\frac{{\rm d}t}{L-\int_{0}^{t/2}\zeta(u){\rm d}u}$
@@
Finally the contribution on $[2-\varepsilon,2]$ is
$\int_{2-\varepsilon}^{2}\Big{(}\ell-\int_{2-\varepsilon}^{t}\zeta(s/2)\,{\rm d}s\Big{)}^{-1}{\rm d}t=\int_{1-\frac{\varepsilon}{2}}^{1}\Big{(}\frac{\ell}{2}-\int_{1-\frac{\varepsilon}{2}}^{v}\zeta(u)\,{\rm d}u\Big{)}^{-1}{\rm d}v=\frac{\varepsilon}{\ell}+O(\varepsilon^{2}).$
Recall that $\ell=\xi^{\prime\prime}(1)^{-1/2}$. Combining terms (and recalling that the second term is negated) yields
@@
Using Corollary 2.5, we now deduce the “no upward outliers” property (1.11) for even models in Proposition 1.2. The idea is that any outlier eigenvalue yields a counterexample to the estimate have we just established.
Proof of Proposition 1.2, Eq. (1.11).
@@
$|{\mathrm{GS}}(H_{N})-{\mathrm{GS}}(\xi)|\leq\delta\quad\quad\text{and}\quad\quad{\mathrm{GS}}_{2,\varepsilon}(H_{N})\leq 2{\mathrm{GS}}(\xi)-\frac{\hat{\zeta}(1)}{2}\left(\sqrt{\xi^{\prime\prime}(1)}-\hat{\zeta}(1)^{-1}\right)^{2}\varepsilon+\iota^{2}\varepsilon.$ (2.6)
@@
(The lower bound on ${\boldsymbol{\lambda}}_{1}$ is trivial since ${\boldsymbol{\lambda}}_{1}(\cdot)\geq{\boldsymbol{\lambda}}_{\lfloor\delta N\rfloor}(\cdot)$.) Let $\theta(\varepsilon)=\arcsin(\sqrt{\varepsilon/2})=\sqrt{\varepsilon/2}+O(\varepsilon)$. With ${\boldsymbol{v}}/\sqrt{N}$ the maximum unit eigenvector of $\nabla_{\mathrm{sp}}^{2}(H_{N}({\boldsymbol{\sigma}}))$, set
@@
Then $R({\boldsymbol{\sigma}}^{+},{\boldsymbol{\sigma}}^{-})=\cos(2\theta(\varepsilon))=1-\varepsilon$. Since $\varepsilon\ll\iota$, Taylor expanding (via Proposition 1.9) gives
@@
However, (2.6) implies the left-hand expression is at most
$-\frac{\hat{\zeta}(1)}{2}\left(\sqrt{\xi^{\prime\prime}(1)}-\hat{\zeta}(1)^{-1}\right)^{2}\varepsilon+\iota^{2}\varepsilon.$
This is a contradiction, completing the proof (with $\iota$ here functioning as $\varepsilon$ in the original statement). ∎
@@ 2.3 No Downward Outliers in Even Models
${\boldsymbol{\xi}}\begin{pmatrix}a_{1,1}&a_{1,2}&a_{1,3}\\ a_{2,1}&a_{2,2}&a_{2,3}\\ a_{3,1}&a_{3,2})&a_{3,3}\end{pmatrix}=\begin{pmatrix}9\xi(a_{1,1})&-3\xi(a_{1,2})&-3\xi(a_{1,3})\\ -3\xi(a_{2,1})&\xi(a_{2,2})&\xi(a_{2,3})\\ -3\xi(a_{3,1})&\xi(a_{3,2})&\xi(a_{3,3})\end{pmatrix}.$
@@
${\sf Tr}\big{(}(AJ_{3}+BJ_{-}+CJ_{*})^{-1}\cdot(DJ_{3}+EJ_{-}+FJ_{*})\big{)}={\sf Tr}\left(\frac{DJ_{3}}{A{\sf Tr}(J_{3})}+\frac{EJ_{-}}{B{\sf Tr}(J_{-}}+\frac{FJ_{*}}{C{\sf Tr}(J_{*})}\right)=\frac{D}{A}+\frac{E}{B}+\frac{F}{C}.$ (2.9)
@@
$\boldsymbol{L}_{3}=LJ_{3}+\frac{\ell J_{-}}{2}+\varepsilon^{2}J_{*}.$
@@
Let $t_{*}=3-4\varepsilon+4\varepsilon^{2}={\sf Tr}(J_{3})$. We take $\Phi_{3}:[0,3]\to{\mathbb{R}}^{3\times 3}$ to be piece-wise linear with $\Phi_{3}(0)=0$, and
@@
It is easy to check that ${\sf Tr}(\Phi_{3}(t))=t$ for all $t\in[0,3]$, and $\Phi_{3}(3)=Q_{3}$. Further, $\Phi$ is again increasing in the positive semi-definite order. This time we set $\alpha_{3}(t)=\zeta\left(\frac{t}{3-t_{*}}\right)$ for $t\in[0,t_{*}]$, and extend $\alpha(t)$ to be constant on $t_{*},3]$. The following three-replica interpolation bound again follows directly from [AZ22b, Theorem 6].
@@
We evaluate the interpolation bound in Proposition 2.6 to first order in $\varepsilon$, writing $\approx$ for equalities up to $O(\varepsilon^{2})$ error. The first term is
$\left\langle{\boldsymbol{\xi}}^{\prime}(Q_{3}),\boldsymbol{L}_{3}\right\rangle\approx L\left\langle{\boldsymbol{\xi}}^{\prime}(Q_{3}),J_{3}\right\rangle$
@@
Recalling the main computation in Proposition 2.2, we find
@@
$\frac{1}{t_{*}}\int_{0}^{t_{*}}\left(L-\left(\int_{0}^{t/t_{*}}\zeta(u)\,{\rm d}u\right)\right)^{-1}{\rm d}t\approx\int_{0}^{1}\hat{\zeta}(v)^{-1}\,{\rm d}v..$
@@
$4\varepsilon\left\langle\left(\boldsymbol{L}_{3}-\int_{0}^{3}\alpha_{3}(s)\Phi_{3}^{\prime}(s)\,{\rm d}s\right),J_{-}/2\right\rangle\approx\frac{4\varepsilon}{\ell}.$
@@
The proof essentially follows from [AMS23a, Theorem 5.1], which deduces transport disorder chaos from shattering. Let us summarize the argument. For $\delta\leq\delta_{*}(\xi)$, Corollary 1.5 provides a decomposition $\{{\mathcal{C}}_{-K},\dots,{\mathcal{C}}_{-1},{\mathcal{C}}_{1},\dots,{\mathcal{C}}_{K}\}$ of $A_{\delta}$ into separated clusters, where without loss of generality ${\mathcal{C}}_{-i}=-{\mathcal{C}}_{i}$ are antipodal pairs (since $\xi$ is even). The clusters ${\mathcal{C}}_{i}$ have small diameter $O(\sqrt{N\delta})$ and are pair-wise separated by a much larger distance $\sqrt{N}/C(\xi)$. Further, Lemma 1.12 implies that
@@
The idea of [AMS23a, Theorem 5.1] is that $\mu_{\beta,\varepsilon_{N}}$ will preserve the clustering (i.e. the property (2.10)), but injects noise into the cluster weights. This is implemented in Proposition 5.4 therein, by proving anti-concentration of the log-ratios $\log\big{(}\mu_{\beta,\varepsilon_{N}}({\mathcal{C}}_{i})/\mu_{\beta,\varepsilon_{N}}({\mathcal{C}}_{j})\big{)}$ for each distinct pair $i\neq j$. The anti-concentration comes from the contribution of ${\boldsymbol{v}}^{\otimes p}$ in $\tilde{H}_{N}$, where $\gamma_{p}>0$ and ${\boldsymbol{v}}$ is chosen such that ±$\pm{\mathcal{C}}_{i},\pm{\mathcal{C}}_{j}$ remain separated when projected along ${\boldsymbol{v}}$. The only input hypothesis that is unavailable in our setting is denoted $\mathsf{S1}$ therein: it is not true that $\mu_{\beta}({\mathcal{C}}_{i})$ is exponentially small for each $i$. However, the proof (in the case of even models) applies with only cosmetic changes as long as ${\mathbb{E}}[\max_{i}\mu_{\beta}({\mathcal{C}}_{i}\cup{\mathcal{C}}_{-i})]\leq 1-C(\xi)^{-1}$, i.e. with uniformly positive probability, the maximum probability of any cluster and its antipodal pair is bounded away from $1$. For this, it is sufficient to show that, if ${\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime}\sim\mu_{\beta}$ are i.i.d. Gibbs samples, then
@@
To show (2.11) we use the “even generic” hypothesis $\sum_{p\in 2{\mathbb{N}}}\frac{1_{\gamma_{p}\neq 0}}{p}=\infty$. It is standard from [Pan13b, Chapter 3.7], see also [Pan16, Section 4], that under this hypothesis, the law of $|R({\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime})|$ (averaged over the disorder $H_{N}$) has an $N\to\infty$ limit $\zeta_{\beta}$ given by the minimizer in the Parisi formula at inverse temperature $\beta$. The precise definition of $\zeta_{\beta}$ is recalled in Section 3, but the only property we need is the well-known fact that $0\in{\rm supp}(\zeta_{\beta})$ whenever $\xi^{\prime}(0)=0$. In particular, (2.11) holds, so the proof of [AMS23a, Theorem 5.1] applies and yields Corollary 1.7. ∎

### v3

@@ 2.2 No Upward Outliers in Even Models
${\mathrm{GS}}_{2,\varepsilon}(H_{N})=\sup_{\begin{subarray}{c}{\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime}\in{\mathcal{S}}_{N}\\ R({\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime})=1-\varepsilon\end{subarray}}\frac{H_{N}({\boldsymbol{\sigma}})+H_{N}({\boldsymbol{\sigma}}^{\prime})}{N}.$ (2.6)
@@
This multi-replica extension of the Parisi formula is given by a similar variational problem as in Proposition 1.1. We choose suitable interpolation parameters to upper bound ${\mathrm{GS}}_{2,\varepsilon}(\xi)$ for small $\varepsilon$, and thus deduce non-existence of outlier eigenvalues at approximate ground states. The general formula replaces $L$ by a $2\times 2$ positive semi-definite matrix $\boldsymbol{L}$, and $\zeta$ by a cumulative distribution function on a monotone path of $2\times 2$ matrices in the positive semi-definite order. The path of matrices is encapsulated by an function $\Phi:[0,2]\to{\mathbb{R}}^{2\times 2}$ with $\Phi(0)=0$ and $\Phi(2)=Q$ and ${\sf Tr}(\Phi(t))=t$ for all $t$, with $\Phi(t)-\Phi(s)$ positive semi-definite for all $t\geq s$. Meanwhile $\alpha:[0,1]\to{\mathbb{R}}_{+}$ is the associated cumulative distribution function, which can be viewed as a positive measure on the range of $\Phi$. For us, the relevant specialization is as follows. Define the $2\times 2$ matrices:
@@
We take as given the minimizing pair $(\zeta,L)=\operatorname*{\rm arg\,min}_{(\zeta,L)\in{\mathcal{K}}}{\mathcal{Q}}(\zeta,L;\xi)$ for the original model. Setting for convenience $\ell=\xi^{\prime\prime}(1)^{-1/2}$, their two-replica generalizations are:
$\boldsymbol{L}=\frac{LJ_{+}+\ell J_{-}}{2};\quad\quad\alpha(t)=\zeta(t/2)/2;\quad\quad\Phi(t)=\begin{cases}\frac{tJ_{+}}{2},\quad\quad\quad\quad\quad\quad\quad\quad\leavevmode\nobreak\ \leavevmode\nobreak\ t\in[0,2-\varepsilon];\\ \frac{(2-\varepsilon)J_{+}}{2}+\frac{{(t-(2-\varepsilon))J_{-}}}{2},\quad t\in[2-\varepsilon,2].\end{cases}$ (2.7)
Note that $\Phi^{\prime}(t)=J_{+}/2$ on $t\in(0,2-\varepsilon)$ and $\Phi^{\prime}(t)=J_{-}/2$ on $t\in(2-\varepsilon,2)$, while $\Phi(1)=Q$ as required.
@@
$2\,{\mathrm{GS}}_{2,\varepsilon}(\xi)$ $\leq\langle\xi^{\prime}(Q),\boldsymbol{L}\rangle-\int_{0}^{2}\Big{\langle}\xi^{\prime\prime}(\Phi(t))\odot\Phi^{\prime}(t),\int_{0}^{t}\alpha(s)\Phi^{\prime}(s)\,{\rm d}s{\Big{\rangle}}\,{\rm d}t$ (2.8)
@@
In evaluating the above expressions, it will help to observe the simple algebraic relations
$J_{+}^{2}=2J_{+},\quad\quad J_{-}^{2}=2J_{-},\quad\quad J_{+}J_{-}=J_{-}J_{+}=\begin{pmatrix}0&0\\ 0&0\end{pmatrix}.$
Since $J_{+}+J_{-}=2I_{2}$ we obtain
$(AJ_{+}+BJ_{-})^{-1}=(A^{-1}J_{+}+B^{-1}J_{-})/4.$
In particular, since $\langle J_{+},J_{+}\rangle=4$ this yields the useful computation:
$\Big{\langle}(AJ_{+}+BJ_{-})^{-1},J_{+}/2\Big{\rangle}=\Big{\langle}(A^{-1}J_{+}+B^{-1}J_{-})/4,J_{+}/2\Big{\rangle}=\frac{1}{2A},\quad\forall A,B\neq 0.$ (2.9)
Similarly we have $\langle J_{-},J_{-}\rangle=4$ and $\Big{\langle}(AJ_{+}+BJ_{-})^{-1},J_{-}/2\Big{\rangle}=\frac{1}{2B}$.
@@
We evaluate the upper bound (2.8) for $2\,{\mathrm{GS}}_{2,\varepsilon}(\xi)$ in Proposition 2.4 to first order in $\varepsilon$. The first term is
$L\big{(}\xi^{\prime}(1)+(1-\varepsilon)\xi^{\prime}(1)\big{)}{+\ell\big{(}\xi^{\prime}(1)-\xi^{\prime}(1-\varepsilon)\big{)}}=2L\xi^{\prime}(1)-L\xi^{\prime\prime}(1)\varepsilon+\ell\xi^{\prime\prime}(1)\varepsilon+O(\varepsilon^{2}).$
@@
$\int\limits_{0}^{2-\varepsilon}\left\langle\frac{\xi^{\prime\prime}(t/2)J_{+}}{2},J_{+}\int_{0}^{t}\zeta(s/2)/4\leavevmode\nobreak\ {\rm d}s\right\rangle{\rm d}t$ ${=\frac{1}{2}\int\limits_{0}^{2-\varepsilon}\xi^{\prime\prime}(t/2)\int_{0}^{t}\zeta(s/2)\leavevmode\nobreak\ {\rm d}s\,{\rm d}t}$
$=2\int\limits_{0}^{1-\frac{\varepsilon}{2}}\xi^{\prime\prime}({v})\left(\int_{0}^{v}\zeta(u)\,{\rm d}u\right){\rm d}v$ (2.10)
$=2\int_{0}^{1}\xi^{\prime\prime}({v})\left(\int_{0}^{v}\zeta(u)\,{\rm d}u\right){\rm d}v-\left(\xi^{\prime\prime}(1)\int_{0}^{1}\zeta(u)\,{\rm d}u\right)\varepsilon+O(\varepsilon^{2}).$
@@
$\Big{\langle}O(\varepsilon)J_{+}+O(1)J_{-},\leavevmode\nobreak\ O(1)J_{+}+O(\varepsilon)J_{-}\Big{\rangle}=O(\varepsilon).$
For the third term, we have $\Phi^{\prime}(t)=J_{+}/2$ on $t\in[0,2-\varepsilon]$. Recalling (2.9), we thus only need to track the $J_{+}$ component in the matrix $\boldsymbol{L}-\int_{0}^{t}\alpha(s)\Phi^{\prime}(s){\rm d}s$ to determine the integral contribution on $[0,2-\varepsilon]$. This yields:
${\frac{1}{2}\int_{0}^{2-\varepsilon}\left(\frac{L}{2}-\frac{1}{2}\int_{0}^{t}\alpha(s)/2\leavevmode\nobreak\ {\rm d}s\right)^{-1}\leavevmode\nobreak\ {\rm d}t}$ $=\int_{0}^{2-\varepsilon}\left(L-\int_{0}^{t/2}\zeta(u){\rm d}u\right)^{-1}\leavevmode\nobreak\ {\rm d}t$
@@
Finally for the contribution on $[2-\varepsilon,2]$ in the third term, we similarly track the $J_{-}$ component in $\boldsymbol{L}-\int_{0}^{t}\alpha(s)\Phi^{\prime}(s){\rm d}s$:
$\frac{1}{2}\int_{2-\varepsilon}^{2}\Big{(}\frac{\ell}{2}-O(\varepsilon)\Big{)}^{-1}{\rm d}t=\frac{\varepsilon}{\ell}+O(\varepsilon^{2}).$
Recall that $\ell=\xi^{\prime\prime}(1)^{-1/2}$, so $\ell\xi^{\prime\prime}(1)\varepsilon+\frac{\varepsilon}{\ell}=2\xi^{\prime\prime}(1)^{1/2}\varepsilon$. Combining the terms above and recalling that the second term contribution from (2.10) is negated in (2.8), we obtain as desired:
@@
Using Corollary 2.5, we now deduce the “no upward outliers” property (1.12) for even models in Theorem 1.3. The idea is that any outlier eigenvalue yields a counterexample to the estimate have we just established.
Proof of Theorem 1.3, Eq. (1.12).
@@
$|{{\mathrm{GS}}_{N}}-{\mathrm{GS}}(\xi)|\leq\delta\quad\quad\text{and}\quad\quad{\mathrm{GS}}_{2,\varepsilon}(H_{N})\leq 2\,{\mathrm{GS}}(\xi)-\frac{\hat{\zeta}(1)}{2}\left(\sqrt{\xi^{\prime\prime}(1)}-\hat{\zeta}(1)^{-1}\right)^{2}\varepsilon+\iota^{2}\varepsilon.$ (2.11)
@@
(The lower bound on ${\boldsymbol{\lambda}}_{1}$ is trivial since ${\boldsymbol{\lambda}}_{1}(\cdot)\geq{\boldsymbol{\lambda}}_{\lfloor\delta N\rfloor}(\cdot)$, so we just focus on the upper bound.) Let
$\theta(\varepsilon)=\arcsin(\sqrt{\varepsilon/2})=\sqrt{\varepsilon/2}+O(\varepsilon).$
With ${\boldsymbol{v}}/\sqrt{N}$ the maximum unit eigenvector of $\nabla_{\mathrm{sp}}^{2}(H_{N}({\boldsymbol{\sigma}}))$, set
@@
Then $R({\boldsymbol{\sigma}}^{+},{\boldsymbol{\sigma}}^{-})=\cos(2\theta(\varepsilon))=1-\varepsilon$. Since $\varepsilon\ll\iota$, Taylor expanding (via (1.24)) gives
@@
However, (2.11) implies the left-hand expression is at most $-\frac{\hat{\zeta}(1)}{2}\big{(}\sqrt{\xi^{\prime\prime}(1)}-\hat{\zeta}(1)^{-1}\big{)}^{2}\varepsilon+\iota^{2}\varepsilon.$ This is a contradiction, completing the proof (with $\iota$ here functioning as $\varepsilon$ in the original statement). ∎
Remark 2.6.
Our motivation for the order parameters in (2.7) was that an optimal choice $({\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime})$ in (2.6) should essentially first decide the midpoint $\frac{{\boldsymbol{\sigma}}+{\boldsymbol{\sigma}}^{\prime}}{2}$, and then choose the increment ${\boldsymbol{\sigma}}-{\boldsymbol{\sigma}}^{\prime}$ to be a top eigenvector of the Hessian evaluated there. For example when near-ground states are not marginally stable, one might expect a positive probability for two low-temperature Gibbs samples for the pair-Hamiltonian (2.6) to agree almost exactly in the first stage, but have orthogonal increments in the second, thus yielding a multiple of $J_{+}$ for the overlap matrix. This suggests that the overlap distribution should be supported on multiples of $J_{+}$ until the second stage. The construction (2.7) according behaves as a concatenation of these two stages. Similar intuitions apply for the $3$-replica case in the next subsection. However while this suggests the general form of the construction, precisely identifying the correct order parameters required some trial and error in each case.
@@ 2.3 No Downward Outliers in Even Models
${\boldsymbol{\xi}}\begin{pmatrix}a_{1,1}&a_{1,2}&a_{1,3}\\ a_{2,1}&a_{2,2}&a_{2,3}\\ a_{3,1}&a_{3,2}&a_{3,3}\end{pmatrix}=\begin{pmatrix}9\xi(a_{1,1})&-3\xi(a_{1,2})&-3\xi(a_{1,3})\\ -3\xi(a_{2,1})&\xi(a_{2,2})&\xi(a_{2,3})\\ -3\xi(a_{3,1})&\xi(a_{3,2})&\xi(a_{3,3})\end{pmatrix}.$
@@
${\sf Tr}\big{(}(AJ_{3}+BJ_{-}+CJ_{*})^{-1}\cdot(DJ_{3}+EJ_{-}+FJ_{*})\big{)}={\sf Tr}\left(\frac{DJ_{3}}{A{\sf Tr}(J_{3})}+\frac{EJ_{-}}{B{\sf Tr}(J_{-}{)}}+\frac{FJ_{*}}{C{\sf Tr}(J_{*})}\right)=\frac{D}{A}+\frac{E}{B}+\frac{F}{C}.$ (2.14)
@@
$\boldsymbol{L}_{3}=LJ_{3}+\frac{\ell J_{-}}{2}+\varepsilon^{2}J_{*}.$ (2.15)
@@
Let $t_{*}=3-4\varepsilon+{2}\varepsilon^{2}={\sf Tr}(J_{3})$. We take $\Phi_{3}:[0,3]\to{\mathbb{R}}^{3\times 3}$ to be piece-wise linear with $\Phi_{3}(0)=0$, and
@@
It is easy to check that ${\sf Tr}(\Phi_{3}(t))=t$ for all $t\in[0,3]$, and $\Phi_{3}(3)=Q_{3}$. Further, $\Phi$ is again increasing in the positive semi-definite order. This time we set $\alpha_{3}(t)=\zeta\left(\frac{t}{{t_{*}}}\right)$ for $t\in[0,t_{*}]$, and extend ${\alpha_{3}}(t)$ to be constant on ${[t_{*},3]}$. The following three-replica interpolation bound again follows directly from [AZ22a, Theorem 3].
@@
We evaluate the interpolation bound in Proposition 2.7 to first order in $\varepsilon$, writing $\approx$ for equalities up to $O(\varepsilon^{2})$ error. For the first term, we expand the definition (2.15) of $\boldsymbol{L}_{3}$ and evaluate the matrix inner-products entry-wise:
$\left\langle{\boldsymbol{\xi}}^{\prime}(Q_{3}),\boldsymbol{L}_{3}\right\rangle\approx L\left\langle{\boldsymbol{\xi}}^{\prime}(Q_{3}),J_{3}\right\rangle{+\frac{\ell}{2}\langle{\boldsymbol{\xi}}^{\prime}(Q_{3}),J_{-}\rangle}$
@@
Recalling the main computation in Proposition 2.2 and Lemma 2.3, we find
@@
${\frac{1}{t_{*}}\int_{0}^{t_{*}}\left(L-\left(\int_{0}^{t}\zeta(s/t_{*})/t_{*}\,{\rm d}s\right)\right)^{-1}{\rm d}t=}\frac{1}{t_{*}}\int_{0}^{t_{*}}\left(L-\left(\int_{0}^{t/t_{*}}\zeta(u)\,{\rm d}u\right)\right)^{-1}{\rm d}t{=}\int_{0}^{1}\hat{\zeta}(v)^{-1}\,{\rm d}v.$
@@
$4\varepsilon\left\langle\left(\boldsymbol{L}_{3}-\int_{0}^{3}\alpha_{3}(s)\Phi_{3}^{\prime}(s)\,{\rm d}s\right)^{{-1}},J_{-}/2\right\rangle\approx\frac{4\varepsilon}{\ell}.$
@@
The proof essentially follows from [AMS25, Theorem 5.1], which deduces transport disorder chaos from shattering. Let us summarize the argument. For $\delta\leq\delta_{*}(\xi)$, Corollary 1.5 provides a decomposition $\{{\mathcal{C}}_{-K},\dots,{\mathcal{C}}_{-1},{\mathcal{C}}_{1},\dots,{\mathcal{C}}_{K}\}$ of $A_{\delta}$ into separated clusters, where without loss of generality ${\mathcal{C}}_{-i}=-{\mathcal{C}}_{i}$ are antipodal pairs since $\xi$ is even. (So far we have not asserted anything about the value of $K$.) The clusters ${\mathcal{C}}_{i}$ have small diameter $O(\sqrt{N\delta})$ and are pair-wise separated by a much larger distance $\sqrt{N}/C(\xi)$. Further, Lemma 1.12 implies that
@@
The idea of [AMS25, Theorem 5.1] is that $\mu_{\beta,\varepsilon_{N}}$ will preserve the clustering (i.e. the property (2.16)), but injects noise into the cluster weights. This is implemented in Proposition 5.4 therein, by proving anti-concentration of the log-ratios $\log\big{(}\mu_{\beta,\varepsilon_{N}}({\mathcal{C}}_{i})/\mu_{\beta,\varepsilon_{N}}({\mathcal{C}}_{j})\big{)}$ for each distinct pair $i\neq j$. The anti-concentration comes from the contribution of ${\boldsymbol{v}}^{\otimes p}$ in $\tilde{H}_{N}$, where $\gamma_{p}>0$ and ${\boldsymbol{v}}$ is chosen such that ±$\pm{\mathcal{C}}_{i},\pm{\mathcal{C}}_{j}$ remain separated when projected along ${\boldsymbol{v}}$. The only input hypothesis that is unavailable in our setting is denoted $\mathsf{S1}$ therein: it is not true that $\mu_{\beta}({\mathcal{C}}_{i})$ is exponentially small for each $i$. However, the proof (in the case of even models) applies with only cosmetic changes as long as ${\mathbb{E}}[\max_{i}\mu_{\beta}({\mathcal{C}}_{i}\cup{\mathcal{C}}_{-i})]\leq 1-C(\xi)^{-1}$, i.e. with uniformly positive probability, the maximum probability of any cluster pair is bounded away from $1$. For this, it is sufficient to show that, if ${\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime}\sim\mu_{\beta}$ are i.i.d. Gibbs samples, then
@@
To show (2.17) we use the “even generic” hypothesis $\sum_{p\in 2{\mathbb{N}}}\frac{1_{\gamma_{p}\neq 0}}{p}=\infty$. It is standard from [Pan13b, Chapter 3.7], see also [Pan16, Section 4], that under this hypothesis, the law of $|R({\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime})|$ (averaged over the disorder $H_{N}$) has an $N\to\infty$ limit $\zeta_{\beta}$ given by the minimizer in the Parisi formula at inverse temperature $\beta$. The precise definition of $\zeta_{\beta}$ is recalled in Section 3, but the only property we need is the well-known fact that $0\in{\rm supp}(\zeta_{\beta})$ whenever $\xi^{\prime}(0)=0$ (recall that ${\rm supp}(\cdot)$ was defined in Proposition 1.13). In particular, (2.17) holds, so the proof of [AMS25, Theorem 5.1] applies and yields Corollary 1.7. ∎